OpenStudy (anonymous):

An article is bought for $7000 and sold 5 years later for $650. Find the depreciation formula, and the value of the article after 2 years. Let y be the value in dollars, and t be time in years.

OpenStudy (anonymous):

@FriedRice

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

hello? @FriedRice

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

itsok

OpenStudy (anonymous):

so you bought a article for $7000

OpenStudy (anonymous):

basically y=_ t+_

OpenStudy (anonymous):

then you sold it after 5 years for $650

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

do you know the formula

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok lets find out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Depreciation = Cost − Salvage Value Life in Number of Periods

OpenStudy (anonymous):

7000-650 /2 6350 /2 3175 = depreciation

OpenStudy (anonymous):

that would be the answer

OpenStudy (anonymous):

y= _ t+ ?

OpenStudy (anonymous):

price after two years?

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

it want the deprecition formula

OpenStudy (anonymous):

@aum

OpenStudy (aum):

y = P(1 - r)^t y is the future value = $650 P is the present value = $70,000 r is the rate of annual depreciation in decimal = ? t is number of years = 5 Plug in and solve for r. Put r in (1), simplify, and that will be the depreciation formula. Then put t = 2 and find y.

OpenStudy (aum):

P is 7000 (not 70,000)

OpenStudy (anonymous):

wait so price after 2 years?

OpenStudy (aum):

You have to first solve for r and find the depreciation formula. Then you can find the price after 2 years.

OpenStudy (anonymous):

whats the equation look like

OpenStudy (aum):

I have given you the equation above.

OpenStudy (aum):

\(\Large y = P(1-r)^t = 7000(1-r)^t\) When t = 5, y = 650. Solve for r.

OpenStudy (anonymous):

13/140

OpenStudy (anonymous):

?

OpenStudy (anonymous):

idk confused

OpenStudy (anonymous):

@aum

OpenStudy (aum):

\(\Large 650 = 7000(1-r)^5 \\ \Large \frac{650}{7000} = (1-r)^5 \\ \Large (1-r)^5 = 0.0928571 \\ \Large 1-r = (0.0928571)^{1/5} = 0.6216745 \\ \Large y = 7000(0.6216745)^t \)

OpenStudy (aum):

That is the depreciation equation. For the second part, put t = 2 and calculate y.

OpenStudy (anonymous):

but y= 7000t?

OpenStudy (aum):

y = 7000(1-r)^t We solved for r and substituted above.

OpenStudy (anonymous):

basically whats the depreciation formula for this eqaution

OpenStudy (anonymous):

it telling me to write it like this y= 7000t + ___

OpenStudy (aum):

Not sure what they are expecting. But y = 7000(0.6216745)^t works nicely because if I put t = 5, y comes out to be ^650 as given in the problem which tells me the equation is correct. But your book may be doing by some other method.

OpenStudy (anonymous):

well no it says it like this _t + _

OpenStudy (aum):

Can you post a screenshot of the question and what it is telling you the equation should look like?

OpenStudy (anonymous):

so 5t+ ???

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (aum):

This is where posting the original question helps. They want a linear depreciation formula. Not the exponential formula that is normally used as I did above. We will redo the problem.

OpenStudy (anonymous):

ok cool

OpenStudy (aum):

y = mt + b when t = 0, y = 7000 7000 = 0 + b. b = 7000 y = mt + 7000 when t = 5, y = 650 650 = m(5) + 7000 650-7000 = 5m -6350 = 5m m = -1270 y = -1270t + 7000 That is your depreciation equation. In the first box enter -1270 and in the second box enter 7000. After 2 years, y = -1270(2) + 7000 y = -2540 + 7000 y = $4,460

OpenStudy (anonymous):

so after 2 years its4,460

OpenStudy (aum):

Yes. But sometimes it may not like the comma or the dollar sign. Try just 4460.

OpenStudy (anonymous):

ok

OpenStudy (aum):

Does it let you know right away if it is correct or not?

OpenStudy (anonymous):

it right :) thanks

OpenStudy (aum):

yw.